# How Performance Pressure Influences AI-Assisted Decision Making

### v2

@@ 1 Introduction
With modern language models facilitating interaction with various AI systems, decision aids are now available across many industries (e.g., medical diagnoses Dilsizian and Siegel, 2014; Duron et al., 2021); financial management Zopounidis and Doumpos, 2002;and criminal recidivism risk McKay, 2020), and when used to complement human abilities, have the potential to outperform either the human or AI working alone. The potential is not necessarily realized, however, because of several challenges: debates on ethical resposibility of decisions Smith (2021); Busuioc (2021); Johnson (2021), the human ability to recognize when AI advice should be taken Schemmer et al. (2023), mental models (biases) regarding AI performance and ability to perform well on subjective tasks Clark et al. (2021); Jones-Jang and Park (2023), and effects of how the AI advice is delivered  Steyvers and Kumar (2023).
@@
Many research directions thus aim to resolve these barriers to complementarity in human-AI performance, measured as appropriate AI advice reliance, which has two dimensions: taking correct AI advice and disregarding incorrect AI advice. Investigations include the effects of showing explanations from explainable AI (XAI) alongside AI system predictions Bansal et al. (2021), introducing cognitive forcing functions when presenting AI advice Buçinca et al. (2021), adjusting AI advice presentation methods Rastogi et al. (2022), and adjusting task framing to account for biases about the types of tasks AI can work with Castelo et al. (2019).
In AI-assisted decision making, the human makes the final decision, bearing full responsibility for its consequences. It is established that performance pressure from responsibility can influence decision making behavior Ashton (1990). But how does it influence AI-assisted decision making?
@@
AI-assisted decision making experiments have considered tasks with stakes that are intrinsically high (loan defaults; Green and Chen, 2019) and low (speed dating; Castelo et al. (2019)), but the stakes have little tangible effect or implication for evaluators. Hence, we observe a gap in the literature of how people rely on AI assistants under performance pressure, that is, when stakes matter personally.
@@
We believe this question is of special significance in the NLP research community, and not only in deployment scenarios. Modern NLP relies on eliciting high-quality data from humans to train models, often with systems in the loop. For example, Dynabench (Kiela et al., 2021) and ANLI (Nie et al., 2020) are datasets where humans work with AI models to create data through finding adversarial or interesting examples. Such datasets are often curated with low personal stakes, e.g., Wadhwa et al. (2024); Krishna et al. (2023, 2024); Lu et al. (2024) and Haduong et al. (2023) crowdsourced annotations and paid hourly wages. Could judiciously applied performance pressure influence the decisions of annotators building research datasets in ways that lead to improvements in data, and by extension AI evaluations and systems?
@@
  • We show how XAI affects advice reliance (both positively and negatively), and how the effects interact with pressure, forming a complex picture about how human behavior changes. These findings suggest opportunities for designing adaptive decision-making environments when different XAI methods are available.
@@ 2 Related Work
It is challenging to design systems that can assist with decision making, because people are influenced by many factors when taking advice, such as their personal expertise Ronayne and Sgroi (2019), the advisor’s reputation Yaniv and Kleinberger (2000), or the style of advice delivery (e.g., inviting or broadcasting; Chhabra et al., 2013; Morrison et al., 2024), resulting in inconsistent advice taking behavior that can be challenging to predict. For example, even if advice is objectively high-quality (e.g., advice based on fact), it may still be discounted in the final decision Wang and Du (2018). AI advisors are expected to complement human decision making and result in higher collaborative performance, compared to individual performance, but recent studies have observed that over- and under-reliance on AI advice result in suboptimal collaboration Bussone et al. (2015); Jacobs et al. (2021).
@@
Algorithmic aversion has been shown to be task-dependent, in line with ideas about how well machines can perform on subjective tasks. When the task is subjective, e.g., predicting speed dating results, Castelo et al. (2019) found increased algorithmic aversion, as opposed to an objective task, e.g., predicting financial outcomes. Hypothesizing that people discount AI advice because they do not trust the AI system, researchers have used explainable AI (XAI) methods and shown the explanations alongside the AI advice. Many studies have observed XAI improving appropriate AI reliance (e.g., Panigutti et al., 2022; Ben David et al., 2021; Lee et al., 2023). Yet Fleiß et al. (2024) observed the opposite: when decisions were about quantifiable skills (e.g., work experience or command of English), rather than soft skills (e.g., diligence or ability to work in teams), adding explanations did not significantly increase AI advice reliance. Jiang et al. (2022) similarly observed XAI failing when the user is too uncertain. Another set of methods aim to mitigate poor AI reliance through cognitive forcing functions—interventions that cause a decision maker to engage in analytical thinking Lambe et al. (2016). For example, Rastogi et al. (2022) successfully employ a cognitive forcing function to reduce anchoring bias—a bias where people weight earlier information higher—by adding a time delay before showing AI advice.
@@
An important environmental factor to consider is the influence of stressors on the human decision maker. Decision making often occurs under time stress or the weight of responsibility, for example. Different stressors can influence decision making in different ways Henderson et al. (2024), and when multiple stressors are present, their compound effect can present itself in additive, synergistic, or antagonistic ways Hale et al. (2017). The influence of stress on AI-assisted decision making is an understudied factor, although in recent work, Swaroop et al. (2024) study how AI-assisted decision makers perform under time pressure, which emerges in real-world settings like operating rooms and search and rescue missions. They study when to provide AI advice in an inherently high-stakes medical diagnosis task, adapted to be approachable to laypeople. Gazit et al. (2023) studied AI-assisted decision making under the pressure of responsibility and observed how responsibility pressures overrode logical reasoning, resulting in lower appropriate AI reliance. The experiment setup involved surveying managers in business organizations, using experts with real responsibilities but asking about their behaviors rather than empirically observing them. Further work is needed to understand the role of responsibility and pressure in AI-assisted decision making.
@@
Performance pressure can be experimentally manipulated through different consequences, e.g., rewards and reputation Stoker et al. (2019). High-quality crowdworker data can be collected by using an appropriate financial incentive in the form of a fair base pay and bonuses. A higher potential reward, or bonus, can increase the pressure on the crowdworker toward higher performance. A common way of presenting the bonus is to frame it as a gain, e.g., “if you do a good job, you can earn a bonus”. Alternatively, the bonus could be framed as a loss, e.g., “if you do a poor job, you will lose your bonus”. The literature in risk aversion, the propensity to play it safe, and loss aversion, the fear of losing out, has observed a stronger pressure effect from framing incentives as a loss rather than a gain Merriman and Deckop (2007). Grgić-Hlača et al. (2022) designed a study investigating how trust in the AI advisor evolves and successfully used the loss framing. In their experiment, users made AI-assisted decisions and updated their mental models of the AI behavior. Here, we are interested in studying the influence of external performance pressure stressors to encourage more appropriate AI reliance (i.e., correctly using AI advice to improve decision-making), rather than studying effects on user trust.
@@ 3 Experiment
H1: We can influence AI advice reliance by manipulating the environmental pressure. Increased performance pressure from monetary incentives framed as a loss will improve appropriate AI advice reliance, and increased pressure from time limits will reduce it.
H2. We can predict the influence of the performance pressure manipulation through measuring the risk aversion level of participants and their trust in the AI advisor. Participants with higher risk aversion will be more careful in their decision making. Participants with higher trust in the AI advisor will have more decisions aligned with the AI advice.
H3. Participants will spend more time considering the AI advice under higher pressure (both monetary and time), regardless of whether they change their decision, because they want to be more careful about their response.
@@
To study how pressure influences AI advice-taking with and without AI explanation aids, we recruit Prolific [^1] https://www.prolific.com/ crowdworkers and task them with judging whether a hotel review is genuine or deceptive. We design a within-subjects experiment manipulating the task environment to apply pressure and run three experimental settings simultaneously, changing the availability of an explanation aid, allowing us to consider all of the above hypotheses at the same time.
@@ 3.1 Dataset
We draw data from the Deceptive Opinion Spam Corpus Ott et al. (2011, 2013), a binary classification task, which contains genuine hotel reviews from travel websites and deceptive reviews written by Amazon Mechanical Turk workers. The task is challenging: human performance is 55%—little better than random chance, ensuring that AI advice taking behavior we observe is not confounded by participants’ prior knowledge or skill.
@@ 3.5 Experiment Setup
[…]
@@
We use a within-subjects design varying the type of pressure (baseline pressure, payment pressure, or time pressure). In the baseline pressure condition, participants are informed they will receive a bonus of 0.5USD for every correct decision. In the payment pressure condition, participants are additionally informed they will lose 0.8USD for every incorrect answer. In the time pressure condition, participants must make a correct decision within 30 seconds to receive the 0.5USD bonus. A timer was displayed to indicate remaining time. If participants ran out of time, the timer would count down negatively. We also use a between-subjects design to study the effects of XAI decision aids (baseline, LIME, or GenAI). The baseline subject group received no explanation for the AI’s prediction, the LIME subject group received a LIME feature importance explanation (where text is highlighted to indicate its association with a label), and the GenAI subject group received a natural language explanation generated by ChatGPT (see §3.2).
@@
The following measures are our dependent variables:
  • Total accuracy of the final answers given by the human.
  • Relative positive AI reliance (RAIR; Appendix A Eq. 1): the ratio of the number of cases where the human relies on AI advice to correct their decision (i.e., they were incorrect before receiving advice and correct after) (Schemmer et al., 2023).
  • Relative positive self-reliance (RSR; Appendix A Eq. 2): the ratio of the number of cases where the human correctly maintains their judgment, disregarding the incorrect AI advice (Schemmer et al., 2023).
For all three measures, higher values are preferable.
@@ 4 Results
After filtering for failed attention checks, we collected responses from 99, 102, and 101 subjects across the three XAI conditions: baseline, LIME, and GenAI, respectively.
@@
Figure 2 summarizes accuracy after receiving AI advice across all conditions. We note that GenAI advice increases accuracy slightly; time pressure slightly decreases accuracy, mirroring Swaroop et al. (2024)’s limited findings of how time pressure influences AI advice-taking.
@@
Figure 1 summarizes the difference in relative positive AI- and self-reliance (respectively, RAIR on the left and RSR on the right), across conditions. We observe varied results with the payment pressure condition. When a natural language explanation is given, RAIR drops by 5% on average; RSR decreases in both no-XAI (6%) and GenAI conditions (8%). Time pressure has the strongest influence, decreasing RAIR and RSR for both XAI methods. Participants completed these tasks faster, averaging 20 seconds per task compared to the baseline and payment conditions (compared to 30 seconds), despite being given 30 seconds to complete each task. Hence we attribute the performance drop in the time pressure condition to rushing.
[…]
@@
Risk aversion scores are not a predictor for AI advice usage, but AI advisor trust is (H2).
[…]
Payment pressure increases time spent considering AI advice (H3).
Although the influence of payment perssure does not have a strong influence on RAIR and RSR, we do observe different decision-making behavior. In the payment pressure condition, participants on average spent 40% longer making the initial decision and 10% longer considering AI advice before making the final decision, as compared to the baseline pressure condition. Despite the increased time spent on tasks, overall performance was largely unaffected, with a minor increase when no XAI is present and a minor decrease when GenAI is present (Figure 2).
@@
In contrast to Schemmer et al. (2023), whose work found that LIME improved RAIR and had a minimal effect on RSR, our results show LIME decreases RAIR compared to no-XAI and reduces RSR in the time pressure condition. Yet LIME synergizes well with payment pressure, improving RAIR over the baseline and maintaining RSR. The RSR change aligns with two participants who stated they relied on the AI significantly and entirely during the time conditions. GenAI explanations improved RAIR and RSR in the baseline condition, but pressure canceled out this effect. However, GenAI marginally improved overall accuracy compared to the other two explanation conditions, indicating the potential for GenAI-style XAI in challenging tasks.

### v3

@@ 1 Introduction
With modern language models facilitating interaction with various AI systems, decision aids are now available across many industries (e.g., medical diagnoses (Dilsizian and Siegel, 2014; Duron et al., 2021); financial management (Zopounidis and Doumpos, 2002); and criminal recidivism risk, McKay, 2020), and when used to complement human abilities, have the potential to outperform either the human or AI working alone. The potential is not necessarily realized, however, because of several challenges: debates on ethical resposibility of decisions (Smith, 2021; Busuioc, 2021; Johnson, 2021), the human ability to recognize when AI advice should be taken (Schemmer et al., 2023), mental models (biases) regarding AI performance and ability to perform well on subjective tasks (Clark et al., 2021; Jones-Jang and Park, 2023), and effects of how the AI advice is delivered  (Steyvers and Kumar, 2023).
@@
Many research directions thus aim to resolve these barriers to complementarity in human-AI performance, measured as appropriate AI advice reliance, which has two dimensions: taking correct AI advice and disregarding incorrect AI advice. Investigations include the effects of showing explanations from explainable AI (XAI) alongside AI system predictions (Bansal et al., 2021), introducing cognitive forcing functions when presenting AI advice (Buçinca et al., 2021), adjusting AI advice presentation methods (Rastogi et al., 2022), and adjusting task framing to account for biases about the types of tasks AI can do (Castelo et al., 2019).
@@
In AI-assisted decision making, the human makes the final decision, bearing full responsibility for its consequences. It is established that performance pressure from responsibility can influence decision making behavior (Ashton, 1990). But how does it influence AI-assisted decision making? AI-assisted decision making experiments have considered tasks with stakes that are intrinsically high (loan defaults; Green and Chen, 2019) and low (speed dating; Castelo et al., 2019), but the stakes have little tangible effect or implication for evaluators. Hence, we observe a gap in the literature of how people rely on AI assistants under performance pressure, that is, when stakes matter personally.
@@
We believe this question is of special significance in the NLP research community, and not only in deployment scenarios. Modern NLP relies on eliciting high-quality data from humans to train models, often with systems in the loop. For example, Dynabench (Kiela et al., 2021) and ANLI (Nie et al., 2020) are datasets where humans work with AI models to create data through finding adversarial or interesting examples. Such datasets are often curated with low personal stakes, e.g., Wadhwa et al. (2024), Krishna et al. (2023, 2024), Lu et al. (2024) and Haduong et al. (2023) crowdsourced annotations and paid hourly wages. Could judiciously applied performance pressure influence the decisions of annotators building research datasets in ways that lead to improvements in data, and by extension AI evaluations and systems?
@@
  • We show how XAI affects advice reliance (both positively and negatively) and interacts with pressure, forming a complex picture about how human behavior changes. These findings suggest opportunities for designing adaptive decision-making environments when different XAI methods are available.
@@ 2 Related Work
Designing systems that can assist with decision making is difficult, because people are influenced by many factors when taking advice, such as their personal expertise (Ronayne and Sgroi, 2019), the advisor’s reputation (Yaniv and Kleinberger, 2000), or the style of advice delivery (e.g., inviting or broadcasting, Chhabra et al., 2013; Morrison et al., 2024), resulting in inconsistent advice taking behavior that can be challenging to predict. For example, even if advice is objectively high-quality (e.g., advice based on fact), it may still be disregarded (Wang and Du, 2018). AI advisors should complement human decision making to achieve higher collaborative performance, compared to individual performance, but recent studies have observed that over- and under-reliance on AI advice result in suboptimal collaboration (Bussone et al., 2015; Jacobs et al., 2021).
@@
Algorithmic aversion has been shown to be task-dependent, in line with ideas about how well machines can perform on subjective tasks. When the task is subjective, e.g., predicting speed dating results, Castelo et al. (2019) found increased algorithmic aversion, as opposed to an objective task, e.g., predicting financial outcomes. Hypothesizing that people discount AI advice because they do not trust the AI system, researchers have used explainable AI (XAI) methods and shown the explanations alongside the AI advice. Many studies have observed XAI positively influencing AI reliance (e.g., Panigutti et al., 2022; Ben David et al., 2021; Lee et al., 2023). Yet Fleiß et al. (2024) observed the opposite: when decisions were about quantifiable skills (e.g., work experience or command of English), rather than soft skills (e.g., diligence or ability to work in teams), adding explanations did not significantly increase AI advice reliance. Jiang et al. (2022) similarly observed XAI failing when the user is too uncertain. Another set of methods aim to mitigate inappropriate AI reliance through cognitive forcing functions—interventions that cause a decision maker to engage in analytical thinking (Lambe et al., 2016). For example, Rastogi et al. (2022) successfully employ a cognitive forcing function to reduce anchoring bias—a bias where people weight earlier information higher—by adding a time delay before showing AI advice.
@@
An important environmental factor to consider is the influence of stressors on the human decision maker. Decision making often occurs under time stress or the weight of responsibility, for example. Different stressors can influence decision making in different ways (Henderson et al., 2024), and when multiple stressors are present, their compound effect can present itself in additive, synergistic, or antagonistic ways (Hale et al., 2017). The influence of stress on AI-assisted decision making is an understudied factor, although in recent work, Swaroop et al. (2024) study how AI-assisted decision makers perform under time pressure, which emerges in real-world settings like operating rooms and search and rescue missions. They study when to provide AI advice in an inherently high-stakes medical diagnosis task, adapted to be approachable to laypeople. Gazit et al. (2023) studied AI-assisted decision making under the pressure of responsibility and observed how responsibility pressures overrode logical reasoning, resulting in lower appropriate AI reliance. The experiment setup involved surveying managers in business organizations, using experts with real responsibilities but asking about their behaviors rather than empirically observing them. Further work is needed to understand the role of responsibility and pressure in AI-assisted decision making.
@@
Performance pressure can be experimentally manipulated through different consequences, e.g., rewards and reputation (Stoker et al., 2019). High-quality crowdworker data can be collected by using an appropriate financial incentive in the form of a fair base pay and bonuses. A higher potential reward, or bonus, can increase the pressure on the crowdworker toward higher performance. A common way of presenting the bonus is to frame it as a gain, e.g., “if you do a good job, you can earn a bonus”. Alternatively, the bonus could be framed as a loss, e.g., “if you do a poor job, you will lose your bonus”. The literature in risk aversion, the propensity to play it safe, and loss aversion, the fear of losing out, has observed a stronger pressure effect from framing incentives as a loss rather than a gain (Merriman and Deckop, 2007). Grgić-Hlača et al. (2022) designed a study investigating how trust in the AI advisor evolves and successfully used the loss framing. In their experiment, users made AI-assisted decisions and updated their mental models of the AI behavior. Here, we are interested in studying the influence of external performance pressure stressors to encourage more appropriate AI reliance (i.e., correctly using AI advice to improve decision-making), rather than studying effects on user trust.
@@ 3 Experiment
H1: We can influence AI advice reliance by manipulating the environmental pressure. Increased performance pressure from monetary incentives framed as a loss will improve appropriate AI advice reliance, and increased pressure from time limits will reduce it (e.g., Zakay, 1993).
H2. The risk aversion level of participants and their trust in the AI advisor can predict the influence of performance pressure (e.g., Essl and Jaussi, 2017). Participants with higher risk aversion will be more careful in their decision making. Participants with higher trust in the AI advisor will have more decisions aligned with the AI advice.
H3. Performance pressure will act as a cognitive forcing function, influencing participants to spend more time when making their decisions, because they want to be more careful about their response.
@@
To study how pressure influences AI advice-taking with and without AI explanation aids, we recruit Prolific [^1] https://www.prolific.com/ crowdworkers and task them with judging whether a hotel review is genuine or deceptive. We design a within-subjects experiment manipulating environmental pressure and run three experimental settings simultaneously, changing the availability of an explanation aid, to consider all of the above hypotheses at the same time.
@@ 3.1 Dataset
We draw data from the Deceptive Opinion Spam Corpus (Ott et al., 2011, 2013), a binary classification task, which contains genuine hotel reviews from travel websites and deceptive reviews written by Amazon Mechanical Turk workers. The task is challenging: human performance is 55%—little better than random chance, ensuring that AI advice taking behavior we observe is not confounded by participants’ prior knowledge or skill.
@@ 3.5 Experiment Setup
We sample 24 reviews from the test set, ensuring a balanced sample of genuine and deceptive reviews, and also of correct vs. incorrect AI predictions. All reviews had positive polarity. We select two additional reviews for practice: one where the AI is correct and one where it is incorrect. The reviews had a length of 45–120 words. Each pressure condition (details below) was assigned a balanced, random assignment of 8 reviews (2 of each {genuine, deceptive} $\times$ {correct, incorrect}), and participants encountered pressure conditions in random order. We included two attention checks and rejected data from participants who failed both.
@@
We use a within-subjects design varying the type of pressure (baseline pressure, payment pressure, or time pressure). In the baseline pressure condition, participants are informed they will receive a bonus of 0.5USD for every correct decision. In the payment pressure condition, participants are additionally informed they will lose 0.8USD for every incorrect answer. In the time pressure condition, participants must make a correct decision within 30 seconds to receive the 0.5USD bonus. A timer was displayed to indicate remaining time. If participants ran out of time, the timer would count down negatively. We also use a between-subjects design to study the effects of XAI decision aids (no-XAI, LIME, or GenAI). The baseline (no-XAI) subject group received no explanation for the AI’s prediction, the LIME subject group received a LIME feature importance explanation (where text is highlighted to indicate its association with a label), and the GenAI subject group received a natural language explanation generated by ChatGPT (see §3.2).
@@
Following Schemmer et al. (2023), we measure AI reliance through relative positive AI reliance (RAIR; Appendix A Eq. 1), defined as the ratio of the number of cases where the human relies on AI advice to correct their decision (i.e., they were incorrect before receiving advice and correct after), and relative positive self-reliance (RSR; Appendix A Eq. 2), defined as the ratio of the number of cases where the human correctly maintains their judgment, disregarding the incorrect AI advice. [^5] An alternative measure frequently used is weight on advice (Kahr et al., 2024; Schmitt et al., 2021; Harvey and Fischer, 1997; Logg et al., 2019; Mucha et al., 2021). We also measure overall task accuracy. For all three measures, higher values are preferable.
@@ 4 Results
After filtering for failed attention checks, we collected responses from 99, 102, and 101 subjects across the three XAI conditions: baseline, LIME, and GenAI, respectively. Details for how significance testing is conducted are in Appendix B.1, and uncorrected p-values are reported in the following section where relevant.
@@
Figure 2 summarizes accuracy after receiving AI advice across all conditions. We note that GenAI advice increases task accuracy significantly over no-XAI ($p<.001$) and LIME ($p<.001$); time pressure slightly decreases accuracy, mirroring Swaroop et al. (2024)’s limited findings of how time pressure influences AI advice-taking.
@@
Figure 1 summarizes the difference in relative positive AI- and self-reliance (respectively, RAIR on the left and RSR on the right), across conditions. We observe varied results with the payment pressure condition. When a natural language explanation is given, RAIR drops by 5% on average; RSR decreases in both no-XAI (6%) and GenAI conditions (8%). Time pressure has the strongest influence: RAIR and RSR decrease in LIME and drop significantly in GenAI ($p<.01$, $p<.01$, respectively). Participants completed these tasks faster, averaging 20 seconds per task compared to the 30 seconds used in the baseline and payment conditions, despite being given 30 seconds to complete each task. Hence we attribute the performance drop in the time pressure condition to rushing.
We build generalized linear mixed models (GLMM) with Gaussian distributions to investigate how pressure and XAI conditions influence observed RAIR and RSR differences (Appendix B.2). All dependent variables are log-transformed, and we consider an effect to be credible if the 97% highest density interval (HDI) excluded zero.
Neither pressure nor XAI support a clear directional effect for RAIR, but both variables are associated with reduced RSR ($\beta=-0.17$, HDI [0.33,0.02][-0.33,-0.02]; $\beta=-0.24$, HDI [0.34,0.07][-0.34,-0.07]). Confirming our observations that time pressure lowers overall accuracy, the models found pressure condition to have a credible negative association with accuracy ($\beta=-0.03$, HDI [.06,.01][-.06,-.01]) .
@@
Risk aversion scores have credible negative associations with RAIR but not RSR. AI advisor trust has overall negative associations with appropriate AI reliance. (H2).
Our GLMM models find both risk aversion measures, MPQ and the Holt-Laury survey, are associated with reduced RAIR ($\beta=-0.33$, HDI [0.61,0.05][-0.61,-0.05]; $\beta=-0.39$, HDI [0.68,0.10][-0.68,-0.10]), but not RSR, despite reporting opposing results: the MPQ personality survey rated participants as risk-loving whereas the Holt-Laury survey rated the same participants as risk-averse, and vice versa. Neither risk score had a credible association with RSR and overall accuracy. We attribute this result to the minimal influence of payment pressure on the dependent variables. Participants’ trust in the AI advisor had a credible negative association with RAIR ($\beta=-0.89$, HDI [1.25,0.43][-1.25,-0.43]) but a positive association with RSR ($\beta=0.74$, HDI $[0.33,1.13]$). Neither risk nor trust measures had reliable associations with overall accuracy. From the postsurvey: 10 participants stated they relied on the AI after realizing the first few judgments aligned with the AI advice, indicating trust in the AI advisor, and 7 participants stated they guessed or followed their gut feeling.
Payment pressure increases decision making time (H3).
We observe different decision-making behavior when users are operating under payment pressure despite its weak influence measured through RAIR and RSR. In the payment pressure condition, participants on average spent 40% longer making the initial decision and 10% longer considering AI advice before making the final decision, as compared to the baseline pressure condition. This suggests payment pressure could act as a cognitive forcing function, encouraging more analysis before making a decision. Despite the increased time spent on tasks, overall performance was largely unaffected, with a minor increase when no XAI is present and a minor decrease when GenAI is available (Figure 2).
@@
In contrast to Schemmer et al. (2023), whose work found that LIME improved RAIR and had a minimal effect on RSR, our results show LIME decreases RAIR compared to no-XAI and reduces RSR in the time pressure condition. Yet LIME synergizes well with payment pressure, improving RAIR over the baseline and maintaining RSR. The RSR change aligns with two participants who stated they relied on the AI significantly and entirely during the time conditions. GenAI explanations improved RAIR and RSR in the baseline condition, but pressure canceled out this effect. However, GenAI marginally improved overall accuracy compared to no-XAI and LIME, indicating the potential for GenAI-style XAI in challenging tasks.